Help


media%2Fc6e%2Fc6ec4629-1766-4bc2-985b-d4Help
Answer #1

a. New sales = 1.10 × $306,000 = $336,600

Increase in profit from new sales = Profit percent × Increase in sales = .15 × ($336,600 – 306,000) = $4,590

Average accounts receivable balance without the discount = Average collection period × Average daily sales = 45 × ($306,000 / 360) = $38,250

Average accounts receivable balance with the discount = Average collection period × Average daily sales = 10 × ( $336,600 / 360) = $9,350

Reduction in accounts receivable = $38,250 – 9,350 = $28,900

The $28,900 cash inflow from reducing accounts receivable will be used to reduce the firm's loan balance.

Interest savings = Interest rate × Loan reduction = .12 × $28,900 = $3,468

Cost of discount = Discount rate × Sales = .02 × $336,600 = $6,732

Net gain (loss) = Increase in profit + Interest savings – Cost of discount = $4,590 + 3,468 – 6,732 = $1,326

b. Yes